OpenStudy (anonymous):

The table shows the low outside temperatures for Monday, Tuesday, and Wednesday. What was the change in the low outside temperature from Tuesday to Wednesday? This function table shows the outside temperature for a given day A. –20.8⁰ B. –12⁰ C. –4.2⁰ D. –3.2⁰ s

OpenStudy (anonymous):

@jacobciezki

OpenStudy (nevermind_justschool):

I need the table :):)

OpenStudy (anonymous):

This morning, Elena wrote a check in the amount of $42.50. This brought her balance down to –$14.25. Suppose n represents Elena's balance before she wrote the check. Which equation can be used to represent this situation? A. n – 14.25 = 42.50 B. n + 42.50 = –14.25 C. n – 42.50 = –14.25 D. n – 14.25 = –42.50

OpenStudy (anonymous):

This morning, Elena wrote a check in the amount of $42.50. This brought her balance down to –$14.25. Suppose n represents Elena's balance before she wrote the check. Which equation can be used to represent this situation? A. n – 14.25 = 42.50 B. n + 42.50 = –14.25 C. n – 42.50 = –14.25 D. n – 14.25 = –42.50